im hving problems with algebra ....need emdiate help ...form 3 student
if x=2,y=3,z=4calculate the vaue of each of the following algebraic expressions :
1. x+5=
2. y+3=
3. z+2= I WOULD APPRICIATE SOME ASSISTANCE IF THIS
SITE IS NOT A SPAM,IM 14 YEARS OLD..THANKS.

Answer by Theo(13342)

you are given that:
x = 2
y = 3
z = 4
first equation is:
x + 5 = ?????
substitute 2 for x to get:
2 + 5 = 7
second equation is:
y + 3 = ?????
substitute 3 for y to get:
3 + 3 = 6
third equation is:
z + 2 = ?????
substitute 4 for z to get:
4 + 2 = 6
all you're doing is substituting the value of the variable for the variable in the equation and then solving the equation.